If a silver spoon weight 56 gms, and has a volume of 5.3cm what is it's density and specific gravity?
1
Expert's answer
2012-05-29T09:47:25-0400
density d = m / V = 56 / 5.3 = 10.57 g/сm^3.
specific gravity = p(sample) /
p(water) = 10.57 / 1 = 10.57.

Answer: 10.57 g/сm^3; 10.57.